# Data-driven Implementations of Various Generalizations of Balanced Truncation

### v2

@@ 4 Practical Considerations in Implementation
$\mathscr{H}(\mu_{i})$ $=\begin{cases}\frac{2\text{Real}(\sigma_{i})}{\mu_{i}-\sigma_{i}}\big[H(\sigma_{i})-H(\mu_{i})\big]H^{*}(\sigma_{i})+H(\mu_{i})D^{T}&\text{if }\sigma_{i}\neq\mu_{i}\\ 2\text{Real}(\sigma_{i})H^{\prime}(\sigma_{i})H^{*}(\sigma_{i})+H(\mu_{i})D^{T}&\text{if }\sigma_{i}=\mu_{i}\end{cases},$ (101)
$t_{w,i}$ $=\Big[R_{\text{X}}+\mathscr{H}^{*}(\mu_{i})\Big]^{-1}R_{\text{X}}^{\frac{1}{2}}.$ (102)
Using these formulas, the algorithms proposed in this paper can be efficiently implemented. Even if samples of $G(s)$ are available along the $j\omega$-axis, one can set a near-zero damping value and still use the above formulas and algorithms without significant loss of accuracy.
5 Numerical Example
@@
This section evaluates the numerical performance of the proposed algorithms using the same $400^{th}$-order RLC circuit model considered in [3]. The MATLAB codes to reproduce the results are provided in [41]. All simulations were performed in MATLAB R2021b on a laptop with a 2 GHz Intel i7 processor and 16 GB of RAM.
Interpolation points for transfer function sampling are computed using formula (78). The imaginary parts $\omega_{i,\sigma}$ of the interpolation points $\sigma_{i}$ consist of 50 logarithmically spaced frequencies between $10^{-1}$ and $10^{4}$. The real parts are determined by setting $\zeta_{i,\sigma}=10^{-5}$. The points $\mu_{i}$ are identical to $\sigma_{i}$. Transfer function samples are obtained numerically from the state-space realization of the RLC model provided in [3]. The quantities $\sqrt{\lambda_{i}(P_{\text{LQG}}Q_{\text{LQG}})}$, $\sqrt{\lambda_{i}(P_{\mathcal{H}_{\infty}}Q_{\mathcal{H}_{\infty}})}$, $\sqrt{\lambda_{i}(P_{\text{PR}}Q_{\text{PR}})}$, $\sqrt{\lambda_{i}(P_{\text{BR}}Q_{\text{BR}})}$, $\sqrt{\lambda_{i}(PQ_{\text{SW}})}$, and $\sqrt{\lambda_{i}(PQ_{\text{S}})}$ are referred to as Hankel-like singular values.
Figures 16 compare the Hankel-like singular values and the relative error $\frac{\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{\infty}}}{\|G(s)\|_{\mathcal{H}_{\infty}}}$. It can be seen that the $25^{th}$-order ROMs generated by intrusive methods (LQG-BT, $\mathcal{H}_{\infty}$-BT, PR-BT, BR-BT, SW-BT, and BST) and their non-intrusive counterparts (DD-LQG-BT, $D$-$\mathcal{H}_{\infty}$-BT, DD-PR-BT, DD-BR-BT, DD-SW-BT, and DD-BST) accurately capture the $20$ most dominant Hankel-like singular values. Moreover, the proposed data-driven algorithms achieve accuracy comparable to that of the intrusive methods for ROMs of orders ranging from $1$ to $25$.
@@
Figure 1: Performance Comparison between LQG-BT and $D$-LQG-BT
@@
Figure 2: Performance Comparison between $\mathcal{H}_{\infty}$-BT and $D$-$\mathcal{H}_{\infty}$-BT
@@
Figure 3: Performance Comparison between PR-BT and $D$-PR-BT
@@
Figure 4: Performance Comparison between BR-BT and $D$-BR-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 5: Performance Comparison between SW-BT and $D$-SW-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 6: Performance Comparison between BST and $D$-BST

### v3

@@ 4.6 Data-driven Projection-based Approximation of BST
$\mathscr{H}(j\nu_{i})$ $=\begin{cases}\dfrac{\epsilon}{2(j\nu_{i}-j\omega_{i})}\big[H(j\omega_{i})-H(j\nu_{i})\big]H^{*}(j\omega_{i})+H(j\nu_{i})D^{T},&\text{if }\omega_{i}\neq\nu_{i},\\[8.61108pt] \dfrac{\epsilon}{2}H^{\prime}(j\omega_{i})H^{*}(j\omega_{i})+H(j\nu_{i})D^{T},&\text{if }\omega_{i}=\nu_{i}.\end{cases}$
5 Numerical Examples
@@
This section evaluates the numerical performance of the proposed algorithms using benchmark dynamical system models for testing MOR algorithms. The MATLAB codes to reproduce the results are provided in [62]. All simulations were performed in MATLAB R2021b on a laptop with a 2 GHz Intel i7 processor and 16 GB of RAM.
Example 1: CD Player
Consider the $120$-th order multiple-input multiple-output (MIMO) model from the benchmark collection of dynamic systems for testing MOR algorithms [10]. For demonstration purposes, the interpolation points consist of two sets of $300$ logarithmically spaced frequencies in $10^{-3},10^{3}]\cup[-10^{3},-10^{-3}]$ rad/sec, forming $150$ conjugate pairs. The right and left interpolation points do not share any common elements. The free parameter $\epsilon$ is set to $10^{-5}$. For QuadBT, quadrature weights are computed for these nodes using the exponential trapezoidal rule [16]. For NI-ADI-BT, the ADI shifts $\epsilon+j\omega_{i}$ are used to approximate $P$, and the ADI shifts $\epsilon+j\nu_{i}$ are used to approximate $Q$. The samples $G(\epsilon+j\omega_{i})$, $G(j\omega_{i})$, $G(\epsilon+j\nu_{i})$, and $G(j\nu_{i})$ are computed numerically using the state-space realization of the CD player. Since the number of samples is moderate, the Sylvester equations and projected Lyapunov equations can be solved directly. Nevertheless, both the exact solutions and their block diagonally dominant approximations are computed and compared. The non-intrusive implementations, which use block diagonally dominant approximations of the projected Lyapunov equations, are marked with an asterisk in the figures. The Hankel singular values of the ROMs of order $25$ produced by BT, QuadBT, NI-ADI-BT, NI-ADI-BT, DD-P-BT, and $D$-P-BT are compared in Fig. 1(a). All non-intrusive BT algorithms capture the $20$ most significant Hankel singular values of the original system. Fig. 1(b) compares the relative error $\frac{\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{\infty}}}{\|G(s)\|_{\mathcal{H}_{\infty}}}$ for ROMs of orders 1251–25. Both intrusive and non-intrusive BT algorithms perform comparably.
(a) Hankel Singular Values Comparison
(b) Relative Error Comparison
Figure 1: Performance Comparison between Intrusive and Non-intrusive BT
Next, the impact of different values of $\epsilon$ on capturing the dominant $25$ Hankel singular values is investigated for $\epsilon=10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$. The relative difference between the exact Hankel singular values and the captured Hankel singular values in the $L_{2}$ norm is shown in Table 2. For $\epsilon=10^{-2}$, $Q_{v}^{-1}$ and $P_{w}^{-1}$ in NI-ADI-BT are ill-conditioned, so the ROM did not capture most Hankel singular values accurately, as indicated by the significant difference in Table 2. Interestingly, when these inverses are replaced with their diagonally dominant approximation, the resulting ROM captured the most significant Hankel singular values accurately, as shown by the small relative difference. For the remaining values of $\epsilon$, no numerical issues arise, leading to accurate ROMs that captured the most significant Hankel singular values well. It can be noted that block-diagonally dominant approximations are promising for avoiding numerical issues and ensuring ROM accuracy. Moreover, these approximations eliminate the need to compute the projected Lyapunov equations. Therefore, we advocate using block-diagonally dominant approximations instead of solving the projected Lyapunov equations.
Table 2: Relative Difference in Hankel Singular values
Method $\epsilon=10^{-2}$ $\epsilon=10^{-3}$ $\epsilon=10^{-4}$ $\epsilon=10^{-5}$
NI-ADI-BT $0.0019$ $8.7849\times 10^{-7}$ $3.8577\times 10^{-7}$ $3.8576\times 10^{-7}$
NI-ADI-BT $3.8613\times 10^{-7}$ $3.8579\times 10^{-7}$ $3.8576\times 10^{-7}$ $3.8576\times 10^{-7}$
$D$-P-BT $3.8576\times 10^{-7}$ $3.8576\times 10^{-7}$ $3.8576\times 10^{-7}$ $3.8576\times 10^{-7}$
$D$-P-BT $3.8576\times 10^{-7}$ $3.8576\times 10^{-7}$ $3.8576\times 10^{-7}$ $3.8576\times 10^{-7}$
Example 2: International Space Station
Consider the $270$-th order MIMO model from the benchmark collection of dynamic systems for testing MOR algorithms [10]. For demonstration purposes, the interpolation points consist of two sets of $500$ logarithmically spaced frequencies in $10^{-3},10^{2}]\cup[-10^{2},-10^{-3}]$ rad/sec, forming $250$ conjugate pairs. The right and left interpolation points do not share any common elements. The free parameter $\epsilon$ is set to $10^{-5}$. For QuadBT, quadrature weights are computed for these nodes using the exponential trapezoidal rule [16]. For NI-ADI-BT, the ADI shifts $\epsilon+j\omega_{i}$ are used to approximate $P$, and the ADI shifts $\epsilon+j\nu_{i}$ are used to approximate $Q$. The samples $G(\epsilon+j\omega_{i})$, $G(j\omega_{i})$, $G(\epsilon+j\nu_{i})$, and $G(j\nu_{i})$ are computed numerically using the state-space realization of the international space station model. Since the number of samples is moderate, the Sylvester equations and projected Lyapunov equations can be solved directly. Nevertheless, both the exact solutions and their block diagonally dominant approximations are computed and compared. The non-intrusive implementations, which use block diagonally dominant approximations of the projected Lyapunov equations, are marked with an asterisk in the figures. The Hankel singular values of the ROMs of order $30$ produced by BT, QuadBT, NI-ADI-BT, NI-ADI-BT, DD-P-BT, and $D$-P-BT are compared in Fig. 2(a). All non-intrusive BT algorithms capture the $20$ most significant Hankel singular values of the original system. Fig. 2(b) compares the relative error $\frac{\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{\infty}}}{\|G(s)\|_{\mathcal{H}_{\infty}}}$ for ROMs of orders 1301–30. As the order of the ROM is increased, the relative error for both intrusive and non-intrusive algorithms declines. Among the non-intrusive methods, QuadBT provides the best performance in this example in terms of capturing the significant Hankel singular values and overall accuracy.
(a) Hankel Singular Values Comparison
(b) Relative Error Comparison
Figure 2: Performance Comparison between Intrusive and Non-intrusive BT
Example 3: Los Angeles Building
Consider the $48$-th order single-input single-output (SISO) model from the benchmark collection of dynamic systems for testing MOR algorithms [10]. For demonstration purposes, the interpolation points consist of two sets of $500$ logarithmically spaced frequencies in $10^{-1},10^{3}]\cup[-10^{3},-10^{-1}]$ rad/sec, forming $250$ conjugate pairs. The right and left interpolation points do not share any common elements. The free parameter $\epsilon$ is set to $10^{-5}$, and the parameter $\gamma$ in $\mathcal{H}_{\infty}$BT is set to $2.5$. For NI-ADI-LQGBT and NI-ADI-$\mathcal{H}_{\infty}$BT, the RADI shifts $\epsilon+j\omega_{i}$ are used to approximate $P_{\mathrm{LQG}}$ and $P_{\mathcal{H}_{\infty}}$, and the shifts $\epsilon+j\nu_{i}$ are used to approximate $Q_{\mathrm{LQG}}$ and $Q_{\mathcal{H}_{\infty}}$. The samples $G(\epsilon+j\omega_{i})$, $G(j\omega_{i})$, $G(\epsilon+j\nu_{i})$, and $G(j\nu_{i})$ are computed numerically using the state-space realization of the building model. Since the number of samples is moderate, the projected Riccati equations can be solved directly. Nevertheless, both the exact solutions and their block diagonally dominant approximations are computed and compared. The non-intrusive implementations that use block diagonally dominant approximations are marked with an asterisk in the figures. The LQG and $\mathcal{H}_{\infty}$ characteristics, defined by $\sqrt{\lambda_{i}(P_{\mathrm{LQG}}Q_{\mathrm{LQG}})}$ and $\sqrt{\lambda_{i}(P_{\mathcal{H}_{\infty}}Q_{\mathcal{H}_{\infty}})}$, respectively, for ROMs of order $25$ produced by LQGBT, NI-ADI-LQGBT, NI-ADI-LQGBT, $D$-P-LQGBT, $D$-P-LQGBT, $\mathcal{H}_{\infty}$BT, NI-ADI-$\mathcal{H}_{\infty}$BT, NI-ADI-$\mathcal{H}_{\infty}$BT, $D$-P-$\mathcal{H}_{\infty}$BT, and $D$-P-$\mathcal{H}_{\infty}$BT are compared in Fig. 3.
(a) LQG Characteristics Comparison
(b) $\mathcal{H}_{\infty}$ Characteristics Comparison
Figure 3: Comparison between Intrusive and Non-intrusive LQGBT and $\mathcal{H}_{\infty}$BT
All non-intrusive LQGBT and $\mathcal{H}_{\infty}$BT algorithms capture the $24$ most significant LQG and $\mathcal{H}_{\infty}$ characteristics of the original system. Fig. 4 compares the relative error $\frac{|G(s)-\hat{G}(s)|-{\mathcal{H}_{\infty}}}{|G(s)|_{\mathcal{H}_{\infty}}}$ for ROMs of orders 1251–25. As the order of the ROM increases, the performance of both intrusive and non-intrusive algorithms becomes comparable.
(a) Relative Error Comparison
(b) Relative Comparison
Figure 4: Accuracy Comparison between Intrusive and Non-intrusive LQGBT and $\mathcal{H}_{\infty}$BT
Example 4: RLC Ladder
Consider the $400$th-order SISO RLC circuit model from [47]. The state-space realization is normalized to ensure $\|G(s)\|_{\mathcal{H}_{\infty}}<1$, making BRBT applicable. The resulting model is stable, positive-real, bounded-real, and minimum-phase, which guarantees the invertibility of $D$, $D^{T}, D+D$$D^{T}$, $I_{m}-D^{T}D$, and $I_{p}-DD^{T}$. Thus, PRBT, BRBT, SWBT, and BST are applicable to this model and are tested in this example. For demonstration purposes, the interpolation points consist of two sets of $200$ logarithmically spaced frequencies in $10^{-1},10^{4}]\cup[-10^{4},-10^{-1}]$ rad/sec, forming $100$ conjugate pairs. The right and left interpolation points do not share any common elements. The free parameter $\epsilon$ is set to $10^{-5}$. For quadrature-based PRBT (QuadPRBT), quadrature-based BRBT (QuadBRBT), and quadrature-based BST (QuadBST), quadrature weights are computed for these nodes using the trapezoidal rule [47]. For NI-ADI-PRBT, NI-ADI-BRBT, NI-ADI-SWBT, and NI-ADI-BST, the ADI shifts $\epsilon+j\omega_{i}$ are used to approximate the generalized controllability Gramians, and the shifts $\epsilon+j\nu_{i}$ are used to approximate the generalized observability Gramians. The samples $G(\epsilon+j\omega_{i})$, $G(j\omega_{i})$, $G(\epsilon+j\nu_{i})$, and $G(j\nu_{i})$ are computed numerically using the state-space realization of the RLC model. Since the number of samples is moderate, the projected Lyapunov and Riccati equations can be solved directly. Nevertheless, both the exact solutions and their block diagonally dominant approximations are computed and compared. The non-intrusive implementations that use block diagonally dominant approximations are marked with an asterisk in the figures. The quantities $\sqrt{\lambda_{i}(P_{\text{PR}}Q_{\text{PR}})}$, $\sqrt{\lambda_{i}(P_{\text{BR}}Q_{\text{BR}})}$, $\sqrt{\lambda_{i}(PQ_{\text{SW}})}$, and $\sqrt{\lambda_{i}(PQ_{\text{S}})}$ are referred to as Hankel-like singular values. Figures 5-8 compare the Hankel-like singular values and the relative error $\frac{\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{\infty}}}{\|G(s)\|_{\mathcal{H}_{\infty}}}$. It can be seen that the $25^{th}$-order ROMs generated by intrusive methods and their non-intrusive counterparts accurately capture the $20$ most dominant Hankel-like singular values. Moreover, the non-intrusive algorithms achieve accuracy comparable to that of the intrusive methods for ROMs of orders ranging from $1$ to $25$.
@@
Figure 5: Performance Comparison between Intrusive and Non-intrusive PRBT
@@
Figure 6: Performance Comparison between Intrusive and Non-intrusive BRBT
@@
Figure 7: Performance Comparison between Intrusive and Non-intrusive SWBT
@@
Figure 8: Performance Comparison between Intrusive and Non-intrusive BST